OpenStudy (realitywillslapyou):

Please help me! I will give you a medal if you do! -2 = 2(3) + b: Solve the equation for the b. Thank you!!

OpenStudy (realitywillslapyou):

Who ever helps me first and the best will get a medal!

OpenStudy (mathstudent55):

-2 = 6 + b -8 = b

OpenStudy (abb0t):

-2 = 6 +b b = -2 - 6 b = ?

OpenStudy (realitywillslapyou):

Thank you mathstudent55!

OpenStudy (realitywillslapyou):

I give to you the medal!